OpenStudy (anonymous):

Find the slope-intercept equation that passes through the (6, 8) and is parallel to the line with the equation 3x + y = 6. A. y = 1/3x + 6 B. y = -3x+ 26 C. y = 3x - 10 D. y = -3x + 30

OpenStudy (anonymous):

so find the slope first.. by putting 3x+y = 6 into linear form.. we see that y = -3x + 6 right? where y = mx+b and m = slope and b = y intercept.

OpenStudy (anonymous):

in a parallel line.. we know that the slope will be the same. so in this case the slope is -3.

OpenStudy (anonymous):

using point slope formula y - y1 = m(x-x1) we plug in our points (6, 8) so y - 8 = m(x - 6)

OpenStudy (anonymous):

since we said the slope was -3 you plug that in and solve for y.

OpenStudy (anonymous):

y = -3(x - 6) + 8 can you solve from there?

OpenStudy (anonymous):

would it be B? y=-3x+26